OpenStudy (anonymous):

Solve the following system. x 2 - y = 3 x - y = -3 The solution set {(-a0, a1), (a2, a3)}.

OpenStudy (lieutenantgeneral):

x 2 - y = 3 = y = x^2 - 3 x - y = -3 = y = x + 3

OpenStudy (anonymous):

i got3,0and 5,0